OpenStudy (anonymous):

Solve x + 7 = 13. –6 6 –20 20

OpenStudy (anonymous):

i got 6 but i need to know if its a posiitve or negative i think that its a positive

OpenStudy (anonymous):

so my answer choice is B

undeadknight26 (undeadknight26):

oi really? its x + 7 = 13 - 7 -7 x = 6

undeadknight26 (undeadknight26):

well gl :D

OpenStudy (anonymous):

oh so its A which is negative 6= -6

OpenStudy (anonymous):

or is it B what my answer was before

undeadknight26 (undeadknight26):

que? no its 6...plug in 6 for x... 6 + 7 = 13 no?

undeadknight26 (undeadknight26):

pm if u have any other questions :D